OpenStudy (anonymous):

A company distributes free boxes of paperclips to all the workers in x offices. Each office has 8 more departments than the number of offices. The number of workers in each department is 4 less than the number of offices. Each worker is given 3 boxes of paperclips. Which expression shows the number of workers in each department? A. x B. x + 8 C. x + 4 D. 3x(x + 8)(x + 4)

OpenStudy (anonymous):

It is x

OpenStudy (texaschic101):

@amistre64 ...can you take a peek at this please

OpenStudy (texaschic101):

@dan815 ...help

OpenStudy (texaschic101):

I keep wanting to say x - 4, but that isn't even a choice

OpenStudy (amistre64):

A company distributes free boxes of paperclips to all the workers in x offices. o = x Each office has 8 more departments than the number of offices. d = o+8 The number of workers in each department is 4 less than the number of offices. w = o - 4

OpenStudy (texaschic101):

so wouldn't workers be x - 4 ?

OpenStudy (amistre64):

thats what i read it as yes .... but its a poorly worded/constructed word problem to start with

OpenStudy (texaschic101):

you did what I did and got d = x + 8 and w = x - 4.....but I didn't know where to go from there

OpenStudy (anonymous):

Just saying right now, D. 3x(x + 8)(x + 4) is incorrect.

OpenStudy (anonymous):

What was the answer lol?